OpenStudy (anonymous):

Tena ate one of the muffins she baked and had 10 left. How many muffins did Tena bake? Write an equation. Then solve the problem.

OpenStudy (anonymous):

X-1=10 x-1+1=10+1 She had 11 muffins before she ate one